If a mixture is composed of 1.3g NaCI and 2.6g CacO3 what percentage of this substance is CaCO3? 0 A 66% B. 50% ° C. 33% D. cannot be determined
Answer #1

percentage of CaCO3 = weight of CaCO3 *100/total weight of compounds ( NaCl+ CaCo3)

                                    =2.6*100/1.3+2.6

                                   = 260/3.9   = 66.66%

A. 66% >>>>> answer